Andy and Ben had a total of $60. If Andy gave been $8.they would have an equal amount.of money. How much money did Ben have in the beginning?
Answer by ankor@dixie-net.com(22740)

Andy and Ben had a total of $60. If Andy gave been $8. they would have an equal amount.of money.
How much money did Ben have in the beginning?
:
write an equation for the original amts
a + b = 60
:
"If Andy gave Ben $8, they would have an equal amount.of money."
a - 8 = b + 8
a - b = 8 + 8
a - b = 16
:
We can use elimination here
a + b = 60
a - b = 16
-------------adding eliminates b, find a
2a = 76
a = 76/2
a = $38, Andy's original amt
then
60 - 38 = $22, Ben's original amt
;
:
You can check these solutions in the statement:
"If Andy gave Ben $8, they would have an equal amount of money. "